dx/1+y = dy/1+x = dz/z


1
Expert's answer
2021-02-19T15:32:06-0500

"\dfrac{dx-dy}{(1+y)-(1+x)} =- \dfrac{d(x-y)}{x-y}"

Also, since "d2=0, \dfrac{dx+dy}{1+y+1+x} = \dfrac{dx+dy+d2}{x+y+2} = \dfrac{d(x+y+2)}{x+y+2}"

Each of the fraction of the given DE is equal to

"-\dfrac{d(x-y)}{x-y}" and "\dfrac{d(x+y+2)}{x+y+2}"

Solving

"-\dfrac{d(x-y)}{x-y} = \dfrac{dz}{z}"

"ln(x-y)^{-1} = lnz + c_1\\
c_1 = \frac{1}{z(x-y)}"

Also, solving

"\dfrac{d(x+y+2)}{x+y+2} = \dfrac{dz}{z}"

"ln(x+y+2) = lnz + c_2\\
c_2 = \frac{x+y+2}{z}"

So the solution is "F(\frac{1}{z(x-y)}, \frac{x+y+2}{z})=0."